A manager has a problem of deciding whether to purchase new equipment and sell it after 5 years or lease the equipment from a vendor for the same period. The price of equipment is $300,000 with insurance and maintenance costs amounting to $11,020 per year. The rent for services is $2,500 per month, payable annually at the end of each year service. What must be the minimum resale value of the equipment in order to justify buying it? Assume that the cost of money is 8% (compounded annually).
Answer #1

Let resale value of the equipment = P

R = 8%

Time = 5 years

Annual lease payment = 2500*12 = $30000

To justify the buying,

PW of the purchase of equipment = PW of the leasing

-300000 - 11020*(1-1/1.08^5)/.08 + P/(1+8%)^5 = -30000*(1-1/1.08^5)/.08

P = (-30000*(1-1/1.08^5)/.08 + 300000 + 11020*(1-1/1.08^5)/.08)*1.08^5

P = $329450.3 or $329450

So, resale value at the end of 5 years will be $329450.3 or $329450.